OpenStudy (anonymous):

The perimeter of a rectangle is 76m. If the width were double and the length were increased by 17 m, the perimeter would be 140 m. What are the length and width of the rectangle?

OpenStudy (saifoo.khan):

HI ABC.

OpenStudy (bahrom7893):

P = 2(a+b) = 76

OpenStudy (bahrom7893):

let a be width and b be length

OpenStudy (bahrom7893):

Actually nevermind let me use w for width and l for length

OpenStudy (bahrom7893):

2(w+b)=76 2(2w+l+17) = 140

OpenStudy (anonymous):

length=23m and width=15m

OpenStudy (bahrom7893):

Solve the two equations

OpenStudy (bahrom7893):

w+b = 38 2w + l + 17 = 70

OpenStudy (saifoo.khan):

2L + 2W = 76 Simplify divide by 2 L + W = 38 or L = 38 - W : "If the width were doubled and the length were increased by 17m, the perimeter would be 120 m." 2(L+17) + 2(2W) = 140 2L + 34 + 4W = 140 2L + 4W = 140 - 34 2L + 4W = 106 Simplify divide by 2, subtract the 1st equation L + 2W = 53 L + W = 38 -------------Subtraction eliminates L, find W W = 15 is the width then L = 38 - 15 L = 23 is the length ; :